# Imaging supermoiré relaxation and conductive domain walls in helical trilayer graphene

### v1

@@ Abstract
In twisted van der Waals materials, local atomic relaxation can significantly alter the underlying electronic structure and properties. Characterizing the lattice reconstruction and the impact of strain is essential to better understand and harness the resulting emergent electronic states. Here, we use a scanning single-electron transistor to image spatial modulations in the electronic structure of helical trilayer graphene, demonstrating relaxation into a superstructure of large domains with uniform moiré periodicity. We further show that the supermoiré domain size is enhanced by strain and can even be altered in subsequent measurements of the same device, while nevertheless maintaining the same local electronic properties within each domain. Finally, we observe higher conductance at the boundaries between domains, consistent with the prediction that they host counter-propagating topological edge modes. Our work confirms that lattice relaxation can produce moiré-periodic order in twisted multilayers, demonstrates strain-engineering as a viable path for designing topological networks at the supermoiré scale, and paves the way to direct imaging of correlation-driven topological phases and boundary modes.
I Introduction
@@
The electronic properties of quantum materials depend sensitively on their lattice structure. Compared to conventional solids, the weak interlayer bonding in van der Waals systems provides additional freedom for atomic-scale relaxation driven by the competing energetics of different stacking configurations and elastic strain [1, 2, 3, 4]. As a result, local twists and strains can develop, enlarging areas with preferred stacking while shrinking disfavored arrangements [5, 6, 7, 8, 9, 10, 11]. This has profound effects on both band structure and the resulting electronic states. For example, different interlayer shifts produce distinct electronic structure [12, 13, 10, 14, 15], and atomic displacements can even alter band topology [16]. Lattice relaxation can also generate new functionality, ranging from alternating ferroelectric domains driven by local symmetry breaking [17, 18, 19, 20, 21] to solitons which host topologically protected boundary modes [22, 23, 24] and can form networks of one-dimensional edge states [25, 5, 6, 7, 26, 27, 8, 9]. Beyond fundamental interest, designing systems where individual edge modes can be controlled may enable applications in topology-based nanoelectronics [22, 23, 25, 28].
@@
Helical trilayer graphene (HTG), in which three graphene layers are consecutively twisted by equal angles , provides an especially rich venue to explore the relationship between lattice relaxation and electronic structure [29, 30, 31, 32, 33, 34, 4, 35, 36, 37, 38, 39]. Nominally, HTG is quasi-periodic, as the moiré pattern formed by the first and second layers is incommensurate with that formed by the second and third layers. However, at small twist angles, HTG is predicted to relax into large triangular domains of moiré-periodic order [32, 33, 34, 4, 35]. This yields two key length scales: a moiré wavelength , which is uniform within the individual domains; and a larger supermoiré wavelength of order , which characterizes the size of the domains. Here is the lattice constant of graphene. HTG has two inequivalent types of domains (denoted as and in Fig. 1a) with different atomic stacking configurations. The domains locally break symmetry, but map onto each other under a transformation (Fig. ED1). Both domains locally host low energy bands that are separated from more dispersive bands by a large band gap.
@@
Here, we conduct single-electron transistor (SET) microscopy of HTG, which reveals periodic modulations of the local electronic structure at the supermoiré scale. Our measurements indicate that HTG relaxes into large domains of moiré-periodic order that locally break symmetry. We demonstrate decoupling between moiré and supermoiré length scales, with strain capable of enlarging and reshaping domains without changing the local electronic character within them. We also observe increased conductivity along the network of domain walls when the domains are gapped, consistent with the prediction of helical edge modes. Our work clarifies the interplay between lattice reconstruction and strain, provides key insight for interpreting transport measurements of mesoscopic devices, and establishes HTG as a promising platform for tunable topological networks.
II Electronic structure and lattice relaxation in HTG
A schematic of the scanning SET measurement setup is illustrated in Fig. 1a (see Methods), and an optical micrograph of the HTG device is shown in Fig. 1b. We plot the inverse electronic compressibility d/d as a function of carrier density at a representative location within the sample in Fig. 1c (red curve). A pair of prominent incompressible peaks occur near cm-2, which we attribute to the superlattice gaps at moiré filling factors . These gaps separate the flat bands near the charge neutrality point from the remote dispersive bands (Fig. 1d). From the carrier densities of the superlattice peaks, we infer a local twist angle (Methods), and as we discuss below, the observation of only one pair of superlattice peaks indicates a single uniform moiré periodicity, despite the presence of two interlayer twist angles. The change in chemical potential across the flat bands is extremely small, meV (Fig. 1c, blue curve). The above characteristics match well to theoretical calculations of the HTG band structure [32, 33, 34, 35, 38].
@@ II Electronic structure and lattice relaxation in HTG
A spatial line cut along the black trajectory in Fig. 1b provides the first indication of lattice relaxation in HTG. The electronic character is uniform (Fig. 1e), and the superlattice peaks disperse only weakly, indicative of small local twist angle variations (Fig. 1f). Importantly, we observe only a single pair of superlattice peaks throughout the spatial line cut. The same observation holds for the entire area of the device (Fig. ED2). This is notable because HTG has two interlayer angles, moiré systems are notoriously plagued by uncontrolled twist angle disorder [45, 46], and spatial variations in each interlayer angle need not be correlated. One would then generically expect two distinct interlayer twist angles in a given location and thus two pairs of superlattice peaks, as observed in other twisted trilayer graphene systems [47, 48]. However, based on the width in carrier density of the superlattice peaks, the interlayer angles would have to match to within throughout the device to produce a single pair of peaks, which we view as extremely unlikely. We also rule out that a large angle mismatch leads to a second superlattice peak outside the accessible range of carrier densities (Supplementary Sec. 2). Instead, we interpret the single pair of peaks as evidence for local relaxation that produces moiré-periodic order within the domains. Theory predicts that this is favored even when starting from two slightly mismatched twist angles, in which case the moiré periodicity is determined by an average of the interlayer angles [32, 33, 34, 4].
III Imaging supermoiré domains
@@
To further demonstrate the effects of lattice relaxation in HTG, we measure the spatial dependence of d/d in the vicnity of the superlattice peak within a 10 area (white box in Fig. 1b). The local twist angle (Fig. 2a) is remarkably homogeneous. In contrast, the peak value of d/d at each location (Fig. 2b) exhibits approximately periodic spatial modulations that are not correlated with features in the twist angle map. The modulations occur on a length scale of a few hundreds of nanometers, significantly exceeding that of the moiré wavelength nm. We therefore ascribe them to variations of the electronic bands in HTG at the supermoiré scale. Similar modulations with the same periodicity are also observed for all other measured incompressible states, both at zero magnetic field (Fig. ED3) and in a perpendicular magnetic field (Fig. ED4).
@@
The locations where d/d at has local minima (dark spots, Fig. 2b) form a triangular lattice. Surrounding these sites, the local maxima form a honeycomb structure (green, Fig. 2b), with narrow lines of lower d/d separating adjacent maxima. This broadly agrees with the theoretical prediction that HTG relaxes into triangular and domains, separated by domain walls where the local stacking order switches (Fig. 1a, right) [32, 33, 34, 4, 35]. In particular, the triangular domains are predicted to be fully gapped at , while their boundaries host gapless modes (see further discussion below). As a result, d/d is maximized at the domain centers, while the gapless modes lead to enhanced compressibility at the domain walls, reaching a local minimum in d/d where they meet at AAA stacking sites (Supplementary Sec. 1). We therefore associate the local minima in Fig. 2b with AAA stacking sites, and the local maxima with the centers of the and domains.
@@
Building on this qualitative understanding, we next assess the length scale of the supermoiré modulation more quantitatively. The experimentally observed distance between adjacent AAA sites is of the same order of magnitude as the supermoiré wavelength nm expected from the local twist angle of the device. However, variations in AAA site separation produce a range of supermoiré domain areas, defined as the triangular areas enclosed by adjacent AAA sites (Fig. 2c). We compare the experimentally observed areas to the theoretically predicted value assuming equal interlayer angles and zero strain. Their ratio is encoded in the color of each triangle in Fig. 2c, which reveals substantial enhancement in the supermoiré domain area relative to the theoretical prediction over a large subset of the device. We emphasize that this cannot be explained by twist angle mismatch because monotonically decreases with increasing (Fig. 2g, Supplementary Sec. 2).
@@
We instead attribute the enhanced supermoiré areas to strain, which has been proposed to modulate domain size in HTG [32] and is ubiquitous in moiré systems [7, 8, 9, 10, 46, 49]. For simplicity, we first discuss isotropic biaxial heterostrain applied to the middle graphene layer, where denotes stretching. The supermoiré wavelength is enhanced within a narrow range centered around a divergence at (Fig. 2h). Strain therefore provides a natural explanation for the increased domain sizes. We note that other strain configurations can also enlarge domains, and uniaxial strain is required to account for the anisotropy of the observed supermoiré domains. We provide a more comprehensive discussion of other strain configurations in Supplementary Sec. 3.
@@
The areal enhancement of the domains, as well as the change in shape, can be attributed to redistribution of strain between measurements. In the new configuration, larger domains are favored to minimize regions associated with higher stacking energy, i.e. domain walls and AAA sites [1, 3, 4, 32], with a relatively smaller cost of increased elastic strain energy. We emphasize that this reshaping of supermoiré domains occurs with almost no change in the local twist angle (i.e. moiré periodicity) or the band structure within the domains, as discussed in more detail at the end of the manuscript. This demonstrates a separation of length scales, providing new opportunities for strain engineering of supermoiré-scale networks without perturbing local physics.
IV Enhanced domain wall conductance
@@
The degree of a.c. enhancement, defined in Fig. 3c, exhibits systematic spatial dependence across the supermoiré domain structure (Fig. 3d). It is largest in the center of the and domains, is reduced at the domain walls, and is smallest at AAA sites. This indicates that the boundaries between domains are more conductive than their interiors (Supplementary Sec. 5). Theoretically, counter-propagating topological edge states with opposite Chern number for the two valleys are predicted along domain walls at in HTG (Fig. 3e). These edge modes manifest as remote Dirac cones within the superlattice gaps of local band structure calculations at the domain wall and AAA site (Fig. 3f-g). While we only probe relative conductivity and thus do not directly address topology, our results are consistent with theoretical predictions.
@@
V Magnetic field dependence
@@
Multiple features in the Landau fan also indicate symmetry breaking, which leads to distinct Hofstadter spectra for the and valleys for . Specifically, the absence of large Hofstadter gaps at high fields near the charge neutrality point can be traced to crossing Hofstadter subbands that originate from different valleys (Fig. ED6, Supplementary Sec. 1). Similarly, the presence of Hofstadter states that extrapolate to intermediate integer are a consequence of the different Hofstadter spectra for each valley. We comment that these can arise without interaction effects. However, the weaker states with odd-integer Chern numbers or zero-field intercepts require spin splitting and likely reflect Hofstadter subband ferromagnetism [50, 51, 52].
@@
In Fig. 4a, we observe neither evidence of moiré-of-moiré physics with states emerging from incommensurate filling factors [53], nor evidence for the presence of two distinct moiré periodicities [47, 48]. This further supports the conclusion that HTG relaxes into large moiré-periodic domains. Moreover, the Landau fan in Fig. 4a is generic across the entire sample (Fig. ED7) and the pattern of Hofstadter states was virtually identical in the second round of measurement (Fig. ED8), despite the significant rearrangement of the supermoiré domain shapes and sizes. These findings emphasize that although heterostrain can modulate the supermoiré pattern, the local electronic properties remain unchanged.
VI Conclusion
@@
In conclusion, we directly imaged spatially modulated electronic states in HTG at supermoiré length scales, demonstrating a profound influence of lattice relaxation on electronic structure. The existence of moiré-periodic domains with conductive edge states at their boundaries has important implications for interpreting mesoscopic transport measurements. Our work strongly motivates future imaging and spectroscopic efforts to study the interactions between edge states and the degree to which they are protected from backscattering [6, 54]. The separation of the moiré and supermoiré length scales indicates that strain can be used to design topological networks, as well as obtain single-domain devices. Due to the ideal quantum geometry of HTG at the magic angle, this may enable the realization of strongly-correlated topological states with quantized response [32, 55, 56].

### v2

# Imaging supermoiré relaxation in helical trilayer graphene

@@ Abstract
In twisted van der Waals materials, local atomic relaxation can alter the underlying electronic structure. Characterizing lattice reconstruction and its susceptibility to strain is essential for understanding emergent electronic states, especially in multilayers where interference between moiré lattices yields larger supermoiré patterns whose energy is highly sensitive to local stacking. Here, we image spatial modulations in the electronic character of helical trilayer graphene which indicate relaxation into a superstructure of large domains with uniform moiré periodicity. We show that the supermoiré domain size is increased by strain and can be altered in the same device while preserving the local properties within each domain. Finally, we observe higher conductance at the domain boundaries, consistent with predictions that they host counter-propagating edge modes. Our work provides real-space visualization of moiré-periodic domains, reveals two independently tunable length scales, and demonstrates strain-engineering as a route toward designing correlated topological networks at the supermoiré scale.
Introduction
@@
The electronic properties of quantum materials depend sensitively on their lattice structure. Compared to conventional solids, the weak interlayer bonding in van der Waals systems provides additional freedom for atomic-scale relaxation driven by the competing energetics of different stacking configurations and elastic strain [27, 2, 54, 26]. As a result, local twists and strains can develop, enlarging areas with preferred stacking while shrinking disfavored arrangements [1, 12, 51, 24, 14, 38, 5]. This has profound effects on both band structure and the resulting electronic states. For example, lattice relaxation in bilayers can generate new functionality, ranging from alternating ferroelectric domains driven by local symmetry breaking [50, 41, 45, 42, 43] to solitons which host topologically protected boundary modes [23, 32, 13] and can form networks of one-dimensional edge states [37, 1, 12, 51, 35, 48, 24, 14]. Additional tunability arises in twisted multilayers, as different interlayer shifts produce distinct electronic structure [15, 30, 38, 16, 10], and interference between moiré patterns leads to a second, larger moiré-of-moiré (supermoiré) length scale which can also modify electronic behavior [47, 5, 11].
@@
Helical trilayer graphene (HTG), in which three graphene layers are consecutively twisted by equal angles , provides an especially rich venue to explore the relationship between lattice relaxation, supermoiré structure, and electronic properties [25, 53, 22, 7, 8, 9, 26, 49, 46, 31, 17, 6]. Nominally, HTG is quasi-periodic, as the moiré pattern formed by the first and second layers is incommensurate with that formed by the second and third layers. However, at small twist angles, HTG is predicted to relax into large triangular domains of moiré-periodic order [7, 8, 9, 26, 49]. This yields a uniform moiré wavelength within the individual domains, whose size is on the order of the supermoiré wavelength . Here is the lattice constant of graphene. Upon reconstruction, HTG has two in-equivalent types of domains (denoted as and in Fig. 1a) with different atomic stacking configurations. The domains locally break symmetry, but map onto each other under a transformation (Extended Data Fig. 1). Both domains locally host low energy bands that are separated from more dispersive bands by a large band gap.
@@
Here, we conduct single-electron transistor (SET) microscopy of HTG, which reveals periodic modulations of the local electronic structure at the supermoiré scale. Our measurements indicate that HTG relaxes into large domains of moiré-periodic order that locally break symmetry. We demonstrate decoupling between moiré and supermoiré length scales, with strain capable of enlarging and reshaping domains without changing the local electronic character within them. We also observe increased conductivity along the network of domain walls when the domains are gapped, consistent with the prediction of helical edge modes. Our work clarifies the interplay between lattice reconstruction and strain, provides key insight for interpreting transport measurements of mesoscopic devices, and establishes that electronic states can be tuned on multiple independent length scales in moiré multilayers.
Electronic structure and lattice relaxation in HTG
A schematic of the scanning SET measurement setup is illustrated in Fig. 1a (see Methods), and an optical micrograph of the HTG device is shown in Fig. 1b. We plot the inverse electronic compressibility d/d as a function of carrier density (Methods) at a representative location within the sample in Fig. 1c (red curve). A pair of prominent incompressible peaks occur near cm-2, which we attribute to the superlattice gaps at moiré filling factors . These gaps separate the flat bands near the charge neutrality point from the remote dispersive bands (Fig. 1d). From the carrier densities of the superlattice peaks, we infer a local twist angle (Methods), and as we discuss below, the observation of only one pair of superlattice peaks indicates a single uniform moiré periodicity, despite the presence of two interlayer twist angles. The change in chemical potential across the flat bands is extremely small, meV (Fig. 1c, blue curve). The above characteristics match well to theoretical calculations of the HTG band structure [7, 8, 9, 49, 17].
@@
A spatial line cut along the black trajectory in Fig. 1b provides the first indication of lattice relaxation in HTG. The electronic character is uniform (Fig. 1e), and the superlattice peaks disperse only weakly, indicative of small local twist angle variations (Fig. 1f). Importantly, we observe only a single pair of superlattice peaks throughout the spatial line cut. The same observation holds for the entire area of the device (Extended Data Fig. 2). This is notable because HTG has two interlayer angles, moiré systems are notoriously plagued by uncontrolled twist angle disorder [40, 19], and spatial variations in each interlayer angle need not be correlated. One would then generically expect two distinct interlayer twist angles in a given location and thus two pairs of superlattice peaks, as observed in other twisted trilayer graphene systems [39, 34]. However, based on the width in carrier density of the superlattice peaks, the interlayer angles would have to match to within throughout the device to produce a single pair of peaks, which we view as extremely unlikely. We also rule out that a large angle mismatch leads to a second superlattice peak outside the accessible range of carrier densities (Supplementary Sec. 2). Instead, we interpret the single pair of peaks as evidence for local relaxation that produces moiré-periodic order within the domains. While we cannot precisely determine the initial two twist angles prior to relaxation, theory predicts that relaxation is favored even when starting from two slightly mismatched angles, in which case the moiré periodicity is determined by an average of the interlayer angles [7, 8, 9, 26].
Imaging supermoiré domains
@@
To further demonstrate the effects of lattice relaxation in HTG, we measure the spatial dependence of d/d in the vicinity of the superlattice peak within a 10 area (white box in Fig. 1b). The local twist angle (Fig. 2a) is remarkably homogeneous. In contrast, the peak value of d/d at each location (Fig. 2b) exhibits approximately periodic spatial modulations that are not correlated with features in the twist angle map. The modulations occur on a length scale of a few hundreds of nanometers, significantly exceeding that of the moiré wavelength nm. We therefore ascribe them to variations of the electronic bands in HTG at the supermoiré scale. Similar modulations with the same periodicity are also observed for all other measured incompressible states, both at zero magnetic field (Extended Data Fig. 3) and in a perpendicular magnetic field (Extended Data Fig. 4).
@@
The locations where d/d at has local minima (dark spots, Fig. 2b) form a triangular lattice. Surrounding these sites, the local maxima form a honeycomb structure (green, Fig. 2b), with narrow lines of lower d/d separating adjacent maxima. This broadly agrees with the theoretical prediction that HTG relaxes into triangular and domains, separated by domain walls where the local stacking order switches (Fig. 1a, right) [7, 8, 9, 26, 49]. In particular, the triangular domains are predicted to be fully gapped at , while their boundaries host gapless modes (see further discussion below). As a result, d/d is maximized at the domain centers, while the gapless modes lead to enhanced compressibility at the domain walls, reaching a local minimum in d/d where they meet at AAA stacking sites (Supplementary Sec. 1). We therefore associate the local minima in Fig. 2b with AAA stacking sites, and the local maxima with the centers of the and domains. We additionally note that the width in density of the superlattice peaks, which is affected by local variations in , show similar modulations at the supermoiré scale. Namely, is maximized at the AAA sites and minimized within the domains (Extended Data Fig. 5). This is consistent with theoretical calculations [7] that show that while the domains of the relaxed HTG lattice structure are moiré-periodic, the AAA sites and domain walls are moiré-aperiodic.
@@
Building on this qualitative understanding, we next assess the length scale of the supermoiré modulation more quantitatively. The experimentally observed distance between adjacent AAA sites is of the same order of magnitude as the supermoiré wavelength nm expected from the local twist angle of the device. However, variations in AAA site separation produce a range of supermoiré domain areas, defined as the triangular areas enclosed by adjacent AAA sites (Fig. 2c). We compare the experimentally observed areas to the theoretically predicted value assuming equal interlayer angles and zero strain. Their ratio is encoded in the color of each triangle in Fig. 2c, which reveals substantial enhancement in the supermoiré domain area relative to the theoretical prediction over a large subset of the device. We emphasize that this cannot be explained by twist angle mismatch because is maximal at (Fig. 2g, Supplementary Sec. 2).
@@
We instead attribute the enhanced supermoiré areas to global heterostrain. We hereafter refer to it only as strain, but note that it is distinct from local lattice relaxation. Such strain has been proposed to modulate domain size in HTG [7] and is ubiquitous in moiré systems [51, 24, 14, 38, 19, 11]. For simplicity, we first discuss isotropic biaxial heterostrain applied to the middle graphene layer, where denotes stretching. The supermoiré wavelength is enhanced within a narrow range centered around a divergence at (Fig. 2h). Heterostrain therefore provides a natural explanation for the increased domain sizes. We note that other strain configurations can also enlarge domains, and uniaxial or shear strain is required to account for the anisotropy of the observed supermoiré domains. We provide a more comprehensive discussion of other strain configurations in Supplementary Sec. 3.
@@
The areal enhancement of the domains, as well as the change in shape, can be attributed to redistribution of strain between measurements. In the new configuration, larger domains are favored to minimize regions associated with higher stacking energy, i.e. domain walls and AAA sites [27, 54, 26, 7], with a relatively small cost of increased elastic energy from lattice relaxation and global heterostrain. We emphasize that this reshaping of supermoiré domains occurs with almost no change in the local twist angle (i.e. moiré periodicity) or the band structure within the domains, as discussed in more detail at the end of the manuscript. Realistic values of heterostrain have little effect on the moiré wavelength, but can parametrically increase the supermoiré scale due to its divergence at moderate strain (Supplementary Sec. 3). Our work demonstrates a separation of length scales, providing new opportunities for strain engineering of supermoiré-scale networks without perturbing local moiré) physics.
Enhanced domain wall conductance
@@
The degree of a.c. enhancement, defined in Fig. 3c, exhibits systematic spatial dependence across the supermoiré domain structure (Fig. 3d). It is largest in the center of the and domains, is reduced at the domain walls, and is smallest at AAA sites. This indicates that the boundaries between domains are more conductive than their interiors (Supplementary Sec. 5). Theoretically, counter-propagating topological edge states with opposite Chern number for the two valleys are predicted along domain walls at in HTG (Fig. 3e). These edge modes manifest as remote Dirac cones within the superlattice gaps of local band structure calculations at the domain wall and AAA site (Fig. 3f-g, see Supplementary Sec. 1). While we only probe relative conductivity and thus do not directly address topology, our results are consistent with theoretical predictions.
@@
Magnetic field dependence
@@
Multiple features in the Landau fan also indicate symmetry breaking, which leads to distinct Hofstadter spectra for the and valleys for . Specifically, the absence of large Hofstadter gaps at high fields near the charge neutrality point can be traced to crossing Hofstadter subbands that originate from different valleys (Extended Data Fig. 7, Supplementary Sec. 1). Similarly, the presence of Hofstadter states that extrapolate to intermediate integer are a consequence of the different Hofstadter spectra for each valley. We comment that these can arise without interaction effects. However, the weaker states with odd-integer Chern numbers or zero-field intercepts require spin splitting and likely reflect Hofstadter subband ferromagnetism [29, 4, 36].
@@
In Fig. 4a, we observe neither evidence of moiré-of-moiré physics with states emerging from incommensurate filling factors [47], nor evidence for the presence of two distinct moiré periodicities [39, 34]. This further supports the conclusion that HTG relaxes into large moiré-periodic domains. Moreover, the Landau fan in Fig. 4a is generic across the entire sample (Extended Data Fig. 8) and the pattern of Hofstadter states was virtually identical in the second round of measurement (Extended Data Fig. 9), despite the significant rearrangement of the supermoiré domain shapes and sizes. These findings emphasize that although heterostrain can modulate the supermoiré pattern, the local electronic properties remain unchanged.
Conclusion
@@
In conclusion, we directly imaged spatially modulated electronic states in HTG at supermoiré length scales, demonstrating a profound influence of lattice relaxation on electronic structure. The existence of moiré-periodic domains with conductive edge states at their boundaries has important implications for interpreting mesoscopic transport measurements. Our work strongly motivates future imaging and spectroscopic efforts to study the interactions between edge states and the degree to which they are protected from backscattering [12, 33]. The observed separation of moiré and supermoiré length scales means that strain can adjust the supermoiré pattern without interrupting the local (moiré) electronic properties. This allows for design of reconfigurable spatially-modulated topological networks, including control over supermoiré symmetry, formation of fractional Chern mosaics [18], and potential for single-domain devices. Due to the ideal quantum geometry of HTG at the magic angle, this may enable the realization of strongly-correlated topological states with quantized response [7, 28, 21]. Lastly, similar effects are likely across a broad range of materials, such as twisted multilayers of transition metal dichalcogenides, 2D magnets, or superconductors, where intertwined topology, correlation, and magnetism can be engineered on the supermoiré scale.
Acknowledgements
We thank Aaron Sharpe, Aviram Uri, Sergio de la Barrera, Li-Qiao Xia, Andrea Young, and Philip Kim for helpful discussions. This work was supported by the QSQM, an Energy Frontier Research Center funded by the U.S. Department of Energy (DOE), Office of Science, Basic Energy Sciences (BES), under Award # DE-SC0021238. K.W. and T.T. acknowledge support from the JSPS KAKENHI (Grant Numbers 21H05233 and 23H02052) and World Premier International Research Center Initiative (WPI), MEXT, Japan. J.C.H. acknowledges support from the Stanford Q-FARM Quantum Science and Engineering Fellowship. Part of this work was performed at the Stanford Nano Shared Facilities (SNSF), supported by the National Science Foundation under award ECCS-2026822.
Author contributions
J.C.H. and Y.L. fabricated the device. J.C.H., Y.L., and Y.H. conducted SET measurements. J.M.M., T.D., Y.L., Y.H., and J.C.H. performed theoretical calculations. T.D. and B.E.F supervised the project. K.W. and T.T provided hBN crystals. All authors contributed to analysis and writing of the manuscript.
Competing interests
The authors declare no competing interest.
Data availability
Data that support the findings in this study are available at https://doi.org/10.5281/zenodo.17365682.
Methods
Device fabrication
The HTG device was fabricated using standard dry transfer techniques with poly (bisphenol A carbonate)/polydimethylsiloxane (PC/PDMS) transfer slides. A monolayer graphene flake was cut into four pieces with a conductive atomic force microscope tip in contact mode. An exfoliated hBN flake (about 50 nm thick) was first used to pick up one of the monolayer graphene flakes, which served as a “sacrificial layer” to prevent subsequent graphene layers from sliding on the hBN surface. We then used the hBN/monolayer graphene stack to sequentially pick the three remaining monolayer graphene flakes at the desired twist angles, with only a small overlap between them and the sacrificial layer to ensure a large bare HTG region remained. We next deposited the hBN/twisted graphene stack on top of a prefabricated bottom gate of few-layer graphite capped by hBN (17 nm), which was annealed in vacuum at 400 ∘C for eight hours to ensure cleanliness of the surface. The full heterostructure was then patterned into a Hall bar geometry with standard electron beam lithography techniques followed by etching and metallization to form edge contacts. The sample at different stages of the fabrication process is shown in Extended Data Fig. 10.
Scanning SET measurements
The SET tips were fabricated by evaporating aluminum onto the apex of a pulled quartz rod. The estimated diameter of each tip at its apex is approximately 100 nm. The tip is then brought to about 50-100 nm above the sample surface. The scanning SET measurements were performed in a Unisoku USM1300 scanning probe microscope system with a microscope head customized for scanning SET operation. All SET measurements are taken at mK or K and are explicitly noted in the figure captions. A bias voltage of 2 mV is applied across the SET tip, and the current response is monitored. A voltage is applied to the graphite bottom gate to tune the carrier density, while a voltage is directly applied to the HTG to minimize tip-induced doping.
The thermodynamic gap of an incompressible state is given by the corresponding step in the chemical potential . is simultaneously and independently measured in two ways with the SET. We obtain d.c. measurements of the chemical potential by tracking the voltage while is maintained at the maximum sensitivity point of the SET tip (i.e., the local electrostatic potential sensed by the tip is held constant). Following the relation , the d.c. chemical potential change can be read out as . Then the change of the chemical potential across the gap is,
where are the densities immediately above and below the gap, respectively.
For the a.c. measurements, a 2-5 mV peak-to-peak a.c. excitation was applied to the sample () and the graphite bottom gate (), at respective frequencies of 823 Hz and 911.7 Hz, unless otherwise stated. The resulting modulations of the SET current and are then measured by lock-in techniques at the corresponding frequencies. The inverse compressibility is obtained by , where is the geometric capacitance between the bottom gate and the sample. The a.c. chemical potential change is obtained by integrating the inverse compressibility. The change of the chemical potential across the gap measured by a.c. is,
where is a small constant background that may need to be subtracted to set where the d.c. chemical potential is flat.
Mathematically, to extract , we fit both and to a logistic function near . The fit takes the form:
where is the fitted change in the chemical potential across an incompressible state, and , , and are other fitting parameters whose specific values are irrelevant to extracting the gap size.
Twist angle determination
The conversion from applied voltages to carrier density is determined by the geometric capacitance between the graphite bottom gate and the HTG sample: , where is the difference in the voltages applied to the bottom gate and the sample. The capacitance between the bottom gate and sample is determined from the slopes of quantum Hall features emerging from charge neutrality in Landau fan measurements, which are quantized according to fundamental constants. This matches well to what is expected from geometrical considerations based on the hBN thickness measured by atomic force microscopy.
The average interlayer twist angle is determined based on the superlattice carrier density , which is found by a Gaussian fit to the superlattice peak in d/d. Here, is the moiré lattice unit cell area and nm is the graphene lattice constant. We note that while we refer to a single angle locally throughout the manuscript, this should be interpreted as the angle corresponding to the single moiré periodicity after relaxation within the HTG domains, and it is not meant to imply that the two interlayer angles are identical before relaxation.